OpenStudy (agar.io):

What is 1/8 of 48? I got 6 @TheSmartOne

TheSmartOne (thesmartone):

@Agar.io @~ You are correct! :) \(\sf\large\frac{48}{8} = \frac{6\times 8}{8} = \boxed{6}\)

OpenStudy (agar.io):

Thanks